# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

### v2

@@ Abstract
Materials and Methods: We utilized data about 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018. Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows. Feature importance and Shapley value analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, notably for African American patients.
Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities.
@@ 1 Introduction
Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 4 or administrative claims data 5. While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression. For instance, Sharma et al. 6 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al. 5 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database. However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics. For example, Tangri et al. 7 used clinical variables like age, gender, eGFR, and albuminuria to predict ESRD progression, and Keane et al. 8 utilized serum creatinine and urine protein levels for high-risk identification. Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices. Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
In the field of explainable AI (XAI) for healthcare, established methods like SHapley Additive exPlanations (SHAP) 9 and Local Interpretable Model-agnostic Explanations (LIME) 10 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models. However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets. At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus. This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making and ultimately improving patient outcomes.
@@ 2 Objective
This research aims to evaluate multiple prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data. Our primary objective is to estimate the probability that a CKD patient will develop ESRD in the future based on the data from a specified observation window prior to the diagnosis of ESRD, represented as follows:
$P\left(\mathrm{ESRD}_{i}=1\mid t,X_{\text{clinical},i}(t),X_{\text{claim},i}(t)\right),$
@@
where $X_{\text{clinical},i}(t)$ and $X_{\text{claim},i}(t)$ denote the clinical and claims data observed up to time $t$ for patient $i$ who progresses to ESRD after the observation window.
We address two key scenarios in our study:
  • Model Robustness: We train and test different analytical models on the same dataset to ensure consistency and comparability, allowing us to select the best-performing model with a focus on robustness and reliability.
  • Clinical Utility: In clinical settings, the likely duration of CKD progression to ESRD is unknown. Our model aims to predict the likelihood of ESRD in the future and suggests an optimal observation window based on the duration of data available on the patient from the time of CKD Stage 3 diagnosis in time steps of 6, 12, 18, 24 and 30 months.
To address these objectives, we focus on identifying the best data-driven observation window while maintaining an acceptable error rate $\epsilon(t)$, represented as follows:
$T_{\text{obs }}=\min\left\{t\mid\epsilon(t)\leq\epsilon^{\prime},t\in\{6,12,18,24,30\}\right\},\vspace{-.5cm}$
where $\epsilon^{\prime}$ is defined as:
$\epsilon^{\prime}=\left\{\begin{array}[]{ll}\textnormal{user-defined error rate}&\textnormal{if specified},\\ \min_{t\in\{6,12,18,24,30\}}\epsilon(t)&\textnormal{otherwise}\end{array}\right.$
@@
Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
3 Materials and Methods
3.1 Data
@@
This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data. The de-identified administrative claims dataset, obtained from a major health insurance provider and approved by their Institutional Review Board, includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. It encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with Chronic Kidney Disease (CKD) and managed by a leading community nephrology practice that shared their de-identified clinical data.
@@
The clinical dataset was sourced from the EHR of the nephrology practice, containing laboratory results, patient demographics, diagnostic details, and medication records from 1998 to 2018. For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data. Usage of the clinical data was approved by the Institutional Review Board of the University.
@@
Preprocessing of the claims data involved removing duplicates, excluding records without any stage of CKD diagnoses, and eliminating entries with negative costs to ensure data integrity. After cleansing, the claims dataset included 5,317,178 individual claims across 7,129 unique patients. The clinical dataset was similarly refined to include 433,421 lab records for 10,326 patients. These datasets were then integrated using unique, de-identified patient IDs, providing a comprehensive view of each patient’s medical and claims history. This integration was crucial for accurately identifying ESRD predictors using multiple analytical models.
3.2 Research Framework
@@
As illustrated in Fig.1, our framework for predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) consists of three key stages: Data Preparation, Modeling & Validation, and Additional Analyses form robustness and clinical utility.
3.2.1 Data Preparation
Data Pre-processing is the foundation of the framework, involving the integration and thorough cleaning of claims and clinical datasets. This Part A stage includes data preprocessing steps such as cleansing, normalization, missing value imputations and log transformations as appropriate for data analysis.
@@ 3.2.1 Data Preparation
Cohort Identification: Following this, in Part B, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period. Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
@@
The study cohort was curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists. We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management. Further refinement excluded patients who progressed to ESRD within the observation window, resulting in a final cohort with uninterrupted, detailed medical and claims histories.
@@
Feature Engineering, depicted in Part C, involved identifying and creating predictive variables essential for modeling. With claims data, we categorized the features used for prediction into two main groups. The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision. These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
@@
The second group, comorbidity-based features, includes patient-specific clinical information such as the duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia. These features are crucial for understanding CKD progression and were statistically analyzed for their significance in predicting ESRD outcomes (as detailed in Table 1(a) and Table 1(b)).
Table 1: Summary of Descriptive and Categorical Statistics for CKD Features
(a) Descriptive statistics for claims-driven and clinical-driven features; means and standard deviations (StD) for ESRD vs. non-ESRD populations. The $p$-value is based on the t-test of the difference between the ESRD and non-ESRD populations. A significance level of $0.05$ is used.
Category Feature Name Overall Mean (StD) ESRD Mean (StD) Non-ESRD Mean (StD) $p$-value
Claims-driven Count of pharmacy claims (CM1) 110 (88) 120 (94) 109 (87) 0.035
Count of inpatient claims (CM2) 3.74 (3.80) 3.75 (3.43) 3.74 (3.87) 0.959
Count of outpatient claims (CM3) 23.69 (22.07) 27.78 (28.75) 23.09 (20.83) 0.003
Count of professional claims (CM4) 89.73 (68.53) 105.37 (77.46) 87.43 (68.01) 0.000
Net cost of pharmacy claims (CM5) 10645 (20303) 12053 (17596) 10440 (20662) 0.110
Net cost of inpatient claims (CM6) 30038 (36080) 33909 (53540) 29440 (32541) 0.121
Net cost of outpatient claims (CM7) 8657 (17495) 9354 (17522) 8554 (17492) 0.415
Net cost of professional claims (CM8) 12137 (13707) 15512 (18657) 11640 (12748) 0.000
Range of claims costs (CM9) 9173 (15910) 11352 (32606) 8852 (11550) 0.147
Standard deviation of claims costs (CM10) 766 (879) 831 (1263) 757 (806) 0.277
Clinical-driven Age at CKD stage 3 first diagnosis (CL1) 72.68 (11) 70.13 (10.37) 73.04 (11) 0.000
CKD stage 3 duration time (days) (CL2) 660 (179) 543 (250) 677 (160) 0.000
Number of emergency department visits (CL3) 2.11 (2.94) 2.18 (2.63) 2.01 (2.99) 0.258
(b) Categorical clinical-driven features; proportions of ESRD vs. non-ESRD populations. The difference between the ESRD and non-ESRD populations is statistically significant for most of the clinical-driven categorical features based on chi-squared test of independence results with a significance level of $0.05$.
Category Feature Name ESRD Proportion Non-ESRD Proportion ${\chi}^{2}$ Statistics $p$-value
Clinical-driven Gender (CL4) (male:female) 60:40 52:48 6.73 0.009
Occurrence of CKD stage 4 (CL5) 43:57 12:88 217.8 0.000
Occurrence of CKD stage 5 (CL6) 5:95 1:99 48.8 0.000
Diabetes (CL7) 73:27 59:41 27.3 0.000
Anemia (CL8) 64:36 62:38 0.58 0.445
Metabolic acidosis (CL9) 25:75 18:82 9.41 0.002
Proteinuria (CL10) 13:87 17:83 3.72 0.054
Secondary hyperparathyroidism (CL11) 33:67 18:82 42.9 0.000
Phosphatemia (CL12) 5:95 3:97 3.55 0.06
Atherosclerosis (CL13) 6:94 14:86 18.04 0.000
Heart failure (CL14) 5:95 11:89 13.7 0.000
Stroke (CL15) 1:99 3:97 4.58 0.032
Conduction & dysrhythmias (CL16) 5:95 16:84 29.9 0.000
Hypertension (CL17) 99:1 97:3 3.1 0.078
@@
  • Demographic Features: These include age, gender, ethnicity and BMI, which have been widely used for understanding CKD progression.
  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, bicarbonate, and albumin, which are important for monitoring CKD. Notably, UACR was not selected due to insufficient readings compared to the other features.
@@
3.2.2 Modeling & Validation
To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns. Our modeling strategy (Fig. 1 Part C) involved partitioning data into training (60%), validation (20%), and testing (20%) sets to ensure model reliability. We optimized performance through hyper-parameter tuning and an iterative process to determine the optimal observation window that balances predictive accuracy and timeliness.
Machine Learning Methods
We employed statistical methods, beginning with logistic regression (LR) 11 as a baseline to estimate the probability of ESRD progression using clinical and claims data features. Beyond LR, we utilized ML techniques such as Random Forest (RF) 12 and Extreme Gradient Boosting (XGBoost) 13. RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting. XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
@@ 3.2.2 Modeling & Validation
Deep Learning Methods
@@
To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 14; 15 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis. This approach captures the temporal evolution of CKD, reflecting its gradual progression and aligning with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data. Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression.
@@
We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 16 for feature extraction, Recurrent Neural Network (RNN) 17, Long Short-Term Memory network (LSTMs 18, Gated Recurrent Unit (GRU) 19 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 20 for sequence modeling. These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision.
Modeling Scenarios: Robustness and Clinical Application
As outlined in the Objective section, our modeling approach addresses two key scenarios:
  • Model Robustness: To ensure model consistency and fair comparability, we train and test on a unified dataset. We address the issue of shrinking cohort size with longer observation windows by selecting cohorts with a minimum of 30 months of data following CKD stage 3 diagnosis. The dataset is partitioned into training (60%), validation (20%), and testing (20%) subsets. By varying training data length from 6 to 30 months within this cohort, we assess model performance under different data availability conditions, thereby evaluating robustness and reliability.
  • Clinical Utility: In clinical practice, the progression timeline from CKD to ESRD is often uncertain. To model this uncertainty, we integrate all time-length cohorts, enabling the prediction of ESRD likelihood and optimal observation windows without precise progression timelines. This mixed-cohort approach enhances model relevance in real-world clinical settings, accommodating variable patient timelines.
3.2.3 Additional Analyses
This stage includes a detailed examination of incorrect predictions, identification of factors contributing to errors, and exploration of potential biases, particularly considering the recent changes to eGFR equations in 2021. These analyses are critical for refining the models and ensuring they deliver fair and accurate predictions.
Analysis of Model Misclassifications
To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set. This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes.

### v3

@@ Abstract
Materials and Methods: We utilized data from 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018. Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows. Feature importance and SHAP analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, especially for African American patients.
Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities for this population.
@@ 1 Introduction
Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 5 or administrative claims data 6; 7. While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression. For instance, Sharma et al. 8 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al. 6 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database. However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics. For example, Tangri et al. 9 used clinical variables such as age, gender, estimated glomerular filtration rate (eGFR)10, and albuminuria to predict ESRD progression, and Keane et al. 11 utilized serum creatinine and urine protein levels for high-risk patient identification. Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices. Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
In the field of explainable AI (XAI) for healthcare, established methods such as SHapley Additive exPlanations (SHAP) 12 and Local Interpretable Model-agnostic Explanations (LIME) 13 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models. However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets. At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus. This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making to ultimately improve patient outcomes.
@@ 2 Objective
This research aims to evaluate prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data. Fig. 1 illustrates the key temporal components of our study design, with the observation window beginning at Stage 3 diagnosis (t=0). During this observation window, patients may progress through different CKD stages (S4, S5); however, we specifically focus on the cohort that does not develop ESRD within the observation window itself. This ensures that our predictive modeling uses only pre-ESRD data to forecast future ESRD development. Our primary objectives are:
@@
  • ESRD Risk Prediction: To estimate the probability that a CKD patient will develop ESRD in the future, based on data from a specified observation window that begins from each patient’s first occurrence of CKD stage 3. This probability is represented as:
    $P\left(\mathrm{ESRD}_{i}=1\mid t>t_{\mathrm{obs}},X_{\text{clinical },i}\left(t_{\mathrm{obs}}\right),X_{\text{claim },i}\left(t_{\mathrm{obs}}\right)\right),$
@@
    where $X_{\text{clinical},i}\left(t_{\text{obs}}\right),X_{\text{claim},i}\left(t_{\text{obs}}\right)$ denote the clinical and claims data observed up to time $t_{\text{obs}}$ for patient $i$.
  • Observation Window Optimization: To identify the optimal observation window $T_{\mathrm{obs}}$ that provides the best performance across all candidate windows $t_{\mathrm{obs}}$ (6, 12, 18, 24, and 30 months), ensuring accurate predictions while minimizing the length of the observation window required:
    $T_{\text{obs }}=\arg\max_{t\in\{6,12,18,24,30\}}\text{ Performance }(t),$
    where Performance $(t)$ is evaluated based on key predictive metrics: F1-score (the harmonic mean of precision and recall) 14, Area Under the Receiver Operating Characteristic curve (AUROC) 15, and Area Under the Precision-Recall Curve (AUPRC) 16.
Figure 1: A timeline illustration of observation windows for CKD progression to ESRD. The observation starts at CKD Stage 3. Objective 1 estimates the probability of ESRD occurrence using clinical and claims data within the observation window, while Objective 2 identifies the optimal observation window that maximizes predictive performance.
@@
This approach balances the trade-off between maximizing predictive accuracy and minimizing the duration of the observation window required for reliable predictions. Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
3 Data Sources
@@
This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data. The de-identified administrative claims dataset was provided by Highmark, a major health insurance provider, and includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. This dataset encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with CKD. The use of this dataset was approved by Highmark’s Institutional Review Board.
@@
The clinical dataset was sourced from the EHR data of TMA, P.C., a leading community nephrology practice. This dataset contains laboratory results, patient demographics, diagnostic details, and medication records spanning from 1998 to 2018. For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data (2009–2018). Usage of the clinical data was approved by the Institutional Review Board of Carnegie Mellon University.
@@
Preprocessing of the claims data involved removing duplicates, excluding records without any CKD stage information, and eliminating entries with negative costs to ensure data integrity. After cleaning, the claims dataset included 5,317,178 individual claims across 7,129 unique patients. The clinical dataset was similarly refined to include 433,421 laboratory records for 10,326 patients. These datasets were then integrated using unique patient IDs to provide a comprehensive view of each patient’s medical and claims history.
4 Methods
As shown in Figure 7, our approach to predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) encompasses three primary stages: Data Preparation, Modeling and Validation, and Additional Analyses to assess robustness and clinical utility.
@@ 4 Methods
4.1 Data Preparation
4.1.1 Data Pre-processing
As mentioned earlier, we performed data cleansing to ensure the integrity of the datasets. This process involved removing duplicate records to prevent redundancy and potential biases in the analysis. We also excluded entries lacking any stage of CKD diagnoses to maintain the datasets’ relevance to our study objectives, and claims records with negative values, which likely indicated data entry errors.
To address missing data issue, we employed Multiple imputation using chained equation (MICE)17, a widely adopted method in healthcare research for its ability to account for multivariable relationships and provide robust imputations. Normalization was applied to ensure that features contributed proportionately to the analysis. We scaled numerical values to a standard range, a step essential for algorithms sensitive to the magnitude of input values. We applied log transformations 18 to variables exhibiting skewed distributions to stabilize variance and approximate normality. This approach is particularly beneficial for handling data with exponential growth patterns, as it mitigates the impact of outliers and enhances the performance of linear models.
In the laboratory data, we observed an additional challenge where many records lacked unit specifications, leading to significant discrepancies in numerical values. To address this, we conducted distribution detection & adjustment19 to identify and correct unit inconsistencies. By analyzing the statistical distribution of each laboratory measurement, we detected shifts indicative of differing units. In consultation with clinical experts, we established expected value ranges for each measurement. Outlier values that deviated markedly from these ranges were adjusted to align with the appropriate units, ensuring consistency and comparability across all records.
4.1.2 Cohort Identification
@@
In Part B of Fig.7, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period. Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
@@
The study cohort was further curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists. We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management, resulting in a final cohort with uninterrupted, detailed clinical and claims histories.
4.1.3 Feature Engineering
@@
Part C depicted in Fig. 7 involves identifying and creating predictive variables essential for modeling. With claims data, we categorized the features used for prediction into two main groups. The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision. These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
@@
The second group, comorbidity-based features, includes patient-specific clinical information such as the patient duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia.
@@
  • Demographic Features: These include age, gender, ethnicity and body mass index (BMI), which have been widely used for understanding CKD progression.
  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, and bicarbonate, which are important for monitoring CKD. Notably, urine albumin-to-creatinine ratio (UACR) 20 and sodium were not selected due to insufficient readings in the dataset.
@@
Following these steps, statistical analyses were performed to compare variables between patients who progressed to ESRD and those who did not. For continuous variables, data were assessed for normality; those meeting normality assumptions were analyzed using independent t-tests to compare group means. For non-normally distributed variables, logarithmic transformations were applied during preprocessing to approximate normality before conducting t-tests. Categorical variables were expressed as frequencies (percentages) and compared using chi-squared tests. Statistical significance was set at $p<0.05$ for all analyses.
4.2 Modeling & Validation
To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns. Our overarching strategy involved partitioning the dataset into training, validation, and testing subsets through stratified sampling, maintaining class distribution across all splits. This approach is particularly crucial in medical research, where class imbalances can significantly impact model performance. To further address the imbalance, we applied SMOTE 21; 22 to the training data, generating synthetic minority samples and enhancing the model’s ability to learn from underrepresented classes.
4.2.1 Machine Learning Methods
We employed statistical methods, beginning with logistic regression (LR) 23 as a baseline model to estimate the probability of ESRD progression using clinical and claims data features. Beyond LR, we utilized ML techniques such as Random Forest (RF) 24 and Extreme Gradient Boosting (XGBoost) 25. We evaluated performance using a standard k-fold cross-validation setup 26. RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting. XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
@@ 4.2.1 Machine Learning Methods
4.2.2 Deep Learning Methods
@@
To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 7; 27 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis. These three-month intervals are indexed as timestamps (timestamp 0 for months 0-3, timestamp 1 for months 3-6, and so on), with a 18-month observation window containing six timestamps in total. This temporal segmentation approach serves multiple purposes: it captures the temporal evolution of CKD, reflects its gradual progression, and aligns with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data. Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression. This timestamp-based structure enables our model to track the progression of disease markers and identify critical temporal patterns that may indicate increased risk of ESRD development.
@@
We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 28 for feature extraction, Recurrent Neural Network (RNN) 29, Long Short-Term Memory network (LSTMs 30, Gated Recurrent Unit (GRU) 31 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 32 for sequence modeling. These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision. We forgo k-fold cross-validation due to high computational costs and training variance from random initialization33. Modern regularization techniques like dropout reduce this need, so we opted for stratified holdout validation to balance efficiency and reliability while maintaining class distributions.
4.2.3 Hyperparameter Tuning
For performance optimization, we employed specialized Python libraries (Python 3.10) to optimize model performance through hyperparameter tuning. For machine learning models, we utilized the H2O framework 34, which provides comprehensive grid search capabilities. The framework systematically explores a predefined set of hyperparameters to identify the optimal combination for model performance. For deep learning models, we leveraged Microsoft’s Neural Network Intelligence (NNI) 35. NNI is an open-source Python toolkit that automates neural architecture search and hyperparameter optimization. It efficiently dispatches and manages trial jobs generated by various tuning algorithms, facilitating the search for optimal neural architectures and hyperparameter configurations.
4.3 Additional Analyses
The comparison of these modeling results helps us identify the optimal observation length and the most reliable and accurate model for predicting CKD progression to ESRD. However, recognizing the limitations of focusing solely on fixed timelines, we extended our analysis to address clinical realities. In practice, the progression timeline from CKD to ESRD is often uncertain when a patient presents for evaluation. To accommodate this variability, we leveraged the selected optimal model and observation window to integrate all time-length cohorts, ranging from 6 to 30 months. This integration enables predictions of ESRD likelihood and optimal observation windows without requiring precise progression timelines, enhancing the model’s relevance and applicability in real-world clinical settings. Building on this framework, we conducted additional analyses to further refine the model, including a detailed investigation of misclassifications and their potential causes, as well as an evaluation of the impact of updated eGFR equations on racial bias in CKD predictions. These efforts aim to ensure fairness and accuracy in the model’s clinical applications.
4.3.1 Analysis of Model Misclassifications